OpenStudy (anonymous):

what is the equation of the line that passes through (1,3) and (-2,-3)

OpenStudy (anonymous):

can you find the slope first? then find the y intercept?

OpenStudy (anonymous):

i suggest you plot the points and connect them with a line

OpenStudy (anonymous):

but you can just use this . slope = (y2 - y1) / (x2 - x1)

OpenStudy (anonymous):

so is the slope 2?

OpenStudy (anonymous):

I don't know how to do any of this that why I needed help.

OpenStudy (anonymous):

yep slope is 2, so have you ever seen this y = mx + b?

OpenStudy (anonymous):

m is the slope b is the y intercept

OpenStudy (anonymous):

so can you plug the slope you found into that equation?

OpenStudy (anonymous):

yea I know but what is the y intercept?

OpenStudy (anonymous):

we are going to find that after you plug in the slope

OpenStudy (anonymous):

so y=2?

OpenStudy (anonymous):

y = 2x + b, 2 = m = slope

OpenStudy (anonymous):

okay what do I do after that?

OpenStudy (anonymous):

ok so we have some points on the line pick one of them to use

OpenStudy (anonymous):

pick either one

OpenStudy (anonymous):

(1,3)

OpenStudy (anonymous):

ok so we have (1, 3) that is the point where x = 1 and y = 3, so lets plug 1 and 3 into this equation y = 2x + b

OpenStudy (anonymous):

so 1=b?

OpenStudy (anonymous):

correct so now you have everything you need for the equation y = 2x + 1

OpenStudy (anonymous):

okay thankyou!

OpenStudy (anonymous):

yw